OpenStudy (anonymous):

What is the approximate length of the lateral surface for the cylinder shown here in net form? A net of a cylinder is shown. It consists of a rectangle and two circles, one above and one below the rectangle. The width of the rectangle is twelve inches. The length of the rectangle is unknown. The diameter of the circle is ten inches. thirty-one point four inches sixty-two point eight inches seventy-eight point five inches three hundred seventy-six point eight inches

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

I don't know. I am sorry.

OpenStudy (anonymous):

Do you know @superhelp101

OpenStudy (superhelp101):

i say thirty-one point four inches

OpenStudy (anonymous):

A certain type of Swiss chocolate comes in a package that's the shape of a triangular prism. If the perimeter of one of the triangular bases is 16 centimeters, what is the surface area of the entire package? A triangular prism is shown. The triangular base has base six centimeters and height four centimeters. The height of the prism is fourteen centimeters. 164 cm2 224 cm2 248 cm2 272 cm2

OpenStudy (anonymous):

@superhelp101

OpenStudy (superhelp101):

1.alright so this the formula SA = wh + lw + lh + ls 2. plug in information. 3. Know the area of two bases = 4 * 6 = 24 4. Area of bottom = 6 * 14= 84 5. Area of two sides = 2 * 5 * 14 = 140 6. Total area = 24+84+140=248

OpenStudy (anonymous):

Thanks! What is the surface area this storage box? A box with length, width, and height of two point five feet each is shown. fifteen point six two five square feet eighteen point seven five square feet thirty-seven point five square feet forty-five square feet

OpenStudy (superhelp101):

It is 18.75 square feet because you have to mult. 2.5 and 2.5 together then x * 3

OpenStudy (anonymous):

A cube has a surface area of one hundred fifty square yards. What is the length of a side? five yards twelve point five yards twenty-five yards thirty-seven point five yards

OpenStudy (anonymous):

@superhelp101

OpenStudy (superhelp101):

5 yards :) * the surface area is 150 square yards.

OpenStudy (anonymous):

Calculate the surface area of this triangular prism. A right triangular prism is shown. The right triangular face has base of twelve meters, height of nine meters, and hypotenuse of fifteen meters. The height of the prism is ten meters. two hundred fifty-five square meters three hundred sixty square meters four hundred sixty-eight square meters four hundred ninety-five square meters

OpenStudy (anonymous):

@superhelp101

OpenStudy (superhelp101):

468m^2 1. What you have to do is calculate the area of the two right triangles so they are congruent 2. Calculi area of one from these and just multiplie it by 2 3. There is two rectangular sides 4. Length 12 & 15 and width 10 so than calculi area for these and there is a rectangular side with width 9 m and length 10 m so you need calculi again area 5. Now u gotta add them up :)

OpenStudy (anonymous):

Caroline's bedroom is the shape of a rectangular prism that measures thirteen feet long, ten feet wide, and 9 feet high. What is the surface area of her room? sixty-four square feet one hundred thirty-four square feet six hundred seventy-four square feet one thousand one hundred seventy square feet

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

@superhelp101